### v3

# COMAE: COMprehensive Attribute Exploration for Zero-shot Hashing

Abstract.
Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios. While considerable success has been achieved, there still exist urgent limitations. Existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes. Also, the continuous-value attributes are not fully harnessed. In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints. By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes. Then COMAE utilizes contrastive learning to comprehensively depict the context of attributes, rather than instance-independent optimization. Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively. Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes. Our codes are available at https://github.com/itsnotacie/TOMM2024-COMAE.
Abstract.
Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios. While considerable success has been achieved, there still exist urgent limitations. Existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes. Also, the continuous-value attributes are not fully harnessed. In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints. By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes. Then COMAE utilizes contrastive learning to comprehensively depict the context of attributes, rather than instance-independent optimization. Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively. Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes. Our source codes are available at https://github.com/YihangZhou0424/NN2024-COMAE.
Prototype Learning, Local Representation, Zero-shot Hashing, Image Zero-shot Hashing, Image Hashing, Image Retrieval.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXjournal: TOMMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Computing methodologies Visual content-based indexing and retrievalccs: Information systems Image search
Image hashing techniques have garnered substantial interest due to their remarkable efficacy in multimedia retrieval and content-based searches for images and documents (Lew et al., 2006; Wang et al., 2017; Ju et al., 2024). Particularly, hashing techniques convert high-dimensional and complicated data into binary codes, which enables efficient searches in the binary Hamming space, thereby contributing to a smaller memory footprint and increased search efficiency (Shi et al., 2022; Li and van Gemert, 2021; Ju et al., 2023). The capability to depict intricate data in a compact binary format not only accelerates retrieval processes but also markedly improves storage efficiency. Consequently, image hashing emerges as a highly valuable asset in situations that necessitate swift and resource-efficient information retrieval.
With the popularity of content and social apps, the growth of Internet images leads to sort of new concepts, presenting a challenge for traditional hashing algorithms that struggle to adapt to unseen data in previously unobserved categories (Venkatesan et al., 2000; Wang et al., 2017; Long et al., 2021). Recent researchers paid attention to the Zero-Shot Hashing (ZSH) issue. Visual attributes, pervasive across prevalent image datasets, serve to delineate discriminative visual properties intrinsic to objects, transcending specific classes and establishing a foundation for addressing the intricacies of previously unseen data (Khan et al., 2018; Chen et al., 2022). Notably, these attributes offer a robust representation for recognizing out-of-distribution samples, making them valuable for capturing similarities in the seen and unseen concepts in Zero-Shot Learning (ZSL) tasks (Arik and Pfister, 2019; Xu et al., 2017). Fig. 1 shows an example of the class-level attributes. Nevertheless, the exploration of visual attributes within the realm of ZSH tasks has been relatively meager. AH (Xu et al., 2017) adopts a straightforward approach by constructing a multi-task architecture to capture the intricate relationships between visual attributes and category classes, thereby facilitating the transfer of knowledge.
While considerable success has been achieved in utilizing visual attributes to establish connections among visual attributes, image features, class labels, and binary hash codes, there still exists urgent limitations (Li and van Gemert, 2021; Xu et al., 2017). Firstly, image locality representations are proven to be critical to zero-shot learning tasks (Li et al., 2021), as the locality exhibits more effective transferability between seen classes and unseen classes. However, existing works ignore the locality representations in deep hashing tasks, which decreases the upper bound of the transferability in hashing algorithms. Secondly, existing works mainly focus on individual image representation learning, optimizing each instance independently. Each dimension of the visual attribute is a continuous value. In contrast to the binary one-hot class encoding, these continuous values encompass richer and more nuanced distance-related information. These informative attributes are aptly poised to capture intricate pairwise relationships, while they are largely ignored by prevailing works. Thirdly, existing works generally employ a multi-task learning framework to independently utilize attribute labels and class labels. Yet their actual relationships are more complicated than such simple presumes. To be specific, two images may belong to the same class but pose disparate attributes. For example, two images with a “bird” class may have “black leg” and “yellow leg”. Similarly, instances with different classes may also exhibit sharing analogous attributes. For example, two images with “bird” and “wapiti” classes may have the same “black leg”. The existing methods may not lead to the global optimum, thus necessitating performing joint optimization to achieve the overall optimum for both tasks.
To tackle the above issues, we conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE. It combines relationships among visual attributes, image features, class labels, and binary hash codes. Specifically, we develop a supervised learning objective that localizes visual attributes for point-wise consistency. By utilizing the attribute prototype network, we aim to capture the fine-grained details and distinctive characteristics of images, enabling more accurate and informative hash codes. To comprehensively harness numerical attribute features, we construct positive and negative instances for each image based on the numerical disparities of attributes and utilize contrastive learning to further model their distance relationships. Thus, our model can capture the subtle variations in attribute values across different images and learn representations that encode the common visual characteristics associated with these attributes. Finally, class-wise optimization is proposed to cohesively learn the hash code, image representation, and visual attributes more effectively, which fosters improved generalization by considering the relationships and dependencies between different classes during the learning process. We jointly learn discriminative global and local features based on the class-level attributes. Extensive experiments on the ZSH benchmark datasets show that COMAE outperforms the state-of-the-art methods, especially in scenarios with a larger number of unseen label classes. It serves to validate the effectiveness of comprehensive attribute explorations.
Figure 1. An example of the class-level attributes.
Our contributions are summarized as follows:
  • We conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE, it depicts the relationships between seen classes and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints.
  • We design an attribute prototype network, a contrastive learning unit, and a joint-optimal label utilization to implement the above three components.
  • Experiments on three popular zero-shot hashing datasets show that COMAE significantly outperforms the state-of-the-art methods, especially in scenarios with a larger number of unseen label classes.
This paper is generally organized as follows. We begin with briefly reviewing related works in Section 2. Then the proposed COMAE is technically introduced in Section 3. We provide extensive experiment results and discussions in Section 4. Finally, we conclude this work in Section 5.
1. Related Work
1.1. Deep Hashing
Image hashing is a popularly used method for approximating nearest neighbor search in large-scale retrieval scenarios (Venkatesan et al., 2000). Shallow image hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008) and ITQ (Gong et al., 2012). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
Over the past decades, deep learning models have shown excellent progress in the field of Computer Vision (Khan et al., 2018; Pourpanah et al., 2022) and Natural Language Processing (Song et al., 2020; Long et al., 2021), many scholars pay attention to leveraging deep neural networks to learn the abstract representations. Deep hashing methods can be achieved through three paths. The first path utilizes limited information, which has a lower upper bound of the capability. ASZH (Shi et al., 2022) is a typical work in the first path, that learns hash functions based on the seeable category labels, for generating binary codes for unseeable data. The second path (Zhang et al., 2019; Guo et al., 2017) generates semantic vectors based on the visual attributes to find a semantic embedding space, then transfers the common knowledge to unseeable classes. TSK, SitNet, and OPZH (Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019) are representative semantic-based works, they conduct a multi-task framework in which the class labels and attribute vectors are co-trained through the hashing network, and the labels are projected into a semantic embedding space. However, there exists an inherent gap between the semantic and visual space, and the alignment of spaces is tough. The third path (Xu et al., 2017) constructs relationships among attributes, image features, binary codes, and class labels. These models are considered to be more generalized when dealing with out-of-distribution and open-set samples (Arik and Pfister, 2019; Gao et al., 2019).
1.2. Zero-shot Learning
Zero-shot learning (ZSL) (Farhadi et al., 2009; Xu et al., 2020; Zheng and Gupta, 2022; Long et al., 2020) is designed to extend object recognition across known and unknown classes via a unified embedding space, where both categories are described by their semantic visual identifiers. Early ZSL methods (Farhadi et al., 2009; Pourpanah et al., 2022) harvest global visual features from networks, whether pre-trained or capable of end-to-end training. Typically, end-to-end frameworks often surpass pre-trained ones by refining visual attributes, thus mitigating biases between the datasets used for training and those for ZSL applications. Nonetheless,such strategies typically produce less than ideal outcomes (Xu et al., 2020), as they struggle to discern the subtle variances between seen and unseen categories. Recent innovations in ZSL predominantly concentrate on methods based on attention (Zheng and Gupta, 2022; Sylvain et al., 2019; Xu et al., 2020), employing visual attributes to identify distinct feature regions, for instance, specific parts of a bird. This emphasis highlights the critical roles of specificity and the arrangement of image features in ZSL. Proposals for prototype networks (Sylvain et al., 2019) strive to pinpoint various attributes within images to bolster the precision of feature localization. Inspired by these insights into visual attributes and the delicate distinctions between recognized and unrecognized classes, our methodology adopts the prototype network to intricately map the relationships between visual attributes and hash codes (Xu et al., 2020).
Figure 2. The architecture of the proposed COMAE. It consists of three modules: a) Point-wise objective aims at improving image locality and attribute representations; b) Pair-wise loss is proposed to learn the representations from individual learning to context-based learning; and c) Class-wise constraint is designed to capture relationships of attributes and class labels.
1.3. Contrastive Learning
Contrastive learning aims at honing representations by juxtaposing similar instance pairs (positive pairs) against dissimilar ones (negative pairs) (Hadsell et al., 2006; Han et al., 2021), enhancing the model’s ability to differentiate between closely related and distinct instances effectively. Many works demonstrate that contrastive learning is able to alleviate the problems of numerous categories and imbalanced class distributions. In addition, it can amplify the subtle distinctions among similar classes, which assists in capturing intricate details that are often overlooked by basic models. (Hadsell et al., 2006) is the first attempt work to improve learned representations by contrasting positive and negative pairs. SimCLR (Chen et al., 2020) utilizes a self-supervised learning architecture that substitutes the conventional memory bank with batch-specific elements, significantly enhancing its performance on the ImageNet dataset. As hash code is a form of representation, recent researches (Yu et al., 2022; Song et al., 2023; Qiu et al., 2021) take attempts to utilize contrastive learning in deep hashing. (Yu et al., 2022) proposes weighted contrastive hashing to explore the fine-grained semantic relations among images. (Song et al., 2023) propose to However, random sampling could result in augmented images with different semantics, deteriorating the performance of contrastive learning.
2. Method
The architecture of the proposed COMAE is shown in Fig. 2. Our proposed COMAE depicts the relationships between seen and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints. To improve the image locality and attribute representations, we propose a point-wise objective. To learn the representations from individual learning to context-based learning, we utilize contrastive learning to build pair-wise loss. For capturing relationships between attributes and class labels, a class-wise constraint is designed.
2.1. Problem Definition
Let $X=\{x_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times H\times W\times C}$ be $N$ images, where $H,W,C$ denote the image height, width, and channel. The matrix $Y=\{y_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C}$ is the label matrix, where $C$ denotes the number of classes. The classes are divided into the seeable class $Y^{s}$ and the unseeable class $Y^{u}$, where $|Y^{s}|\cup|Y^{u}|=|C|,|Y^{s}|\cap|Y^{u}|=\emptyset$. Let $A=\{a_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times K}$ denote the class-level attribute vector in $K$ dimensions. By using training samples which contain labeled images and attributes from seeable classes, i.e., $S=\{x,y,a|x\in X,y\in Y^{s},a\in A\}$, we aim to learn a hashing model $\mathcal{H}\colon X\rightarrow b_{i}\in\{-1,1\}^{l}$, where $b_{i}$ represents the learned $l$-bit binary hash codes. In the testing process, attribute vectors of unseen classes are also known. In this paper, zero-shot hashing focuses on predicting the hashing label of images from unseen classes, i.e., $X\rightarrow Y^{u}$.
2.2. Point-wise Attribute Consistency
Image locality representations are proven to be practical in transferring knowledge from seen to unseen classes (Xu et al., 2020; Li et al., 2021). For more subtly depicted image and attribute characterizations, we first conduct point-wise consistency constraint, which takes the attribute regression loss as the supervision. Prototype learning has proven to be robust in dealing with open-set recognition and out-of-distribution samples (Xu et al., 2020; Tan et al., 2022) in recent years. In the prototype learning scheme, each category is represented by a prototype. This prototype can be the mean of the sample in the category, the center of mass, or some other representative sample. Thus we introduce a prototype module aiming at capturing the relationships of local features and visual attributes.
Let $f_{i,j}(x)\in\mathbb{R}^{C}$ denote the image representation at spatial position (or location) $(i,j)$. It depicts the information on local image regions. Then we learn sorts of attribute prototypes $P=\{p_{k}\in\mathbb{R}^{C}\}_{k=1}^{K}$ to predict attributes from the local visual characterizations. Let $p_{k}$ denote the learnable prototype embeddings for the $k$-th attribute. As shown in Fig. 2, $p_{1}$ and $p_{2}$ correspond to the prototypes for “red beak” and “yellow leg” respectively. For each attribute, we produce a similarity map $M^{k}\in\mathbb{R}^{H\times W}$, where each element is computed by a non-linear network $\Theta(\cdot)$ between for $p_{k}$ and each local feature. To associate each visual attribute with its closest local feature, we gain the predicted $k$-th attribute $\hat{a_{k}}$ by taking the maximum value in the similarity map $M^{k}$,
(1) $M_{i,j}^{k}$ $=\Theta\left(p_{k},f_{i,j}(x)\right),$
$\hat{a}_{k}$ $=\max_{i,j}M_{i,j}^{k}.$
This prototype captures the common characteristics of the category and plays a key role in the expression of the entire category. The visual attribute vectors supervise the learning of attribute prototypes. We formalize the attribute prediction task as a regression problem and then minimize the Mean Square Error (MSE) between the ground truth attributes and the predicted attribute score $\hat{a}$. By optimizing the regression objective, we enforce the local features to encode pivotal semantic attributes, improving the locality of representations, and further improving the quality of hashing codes,
(2) $\mathcal{L}_{\text{pointwise}}=\sum_{i}^{N}\frac{1}{N}||a_{i}-\hat{a_{i}}||^{2}.$
Furthermore, we refine the similarity map to intensively focus on its peak region. To achieve this, we employ a compactness regularizer (Zheng et al., 2017) on each similarity map $M^{k}$, ensuring a concentrated similarity assessment. The equation is defined as,
(3) $\mathcal{L}_{\text{reg}}=\sum^{K}_{k}\sum^{H}_{h}\sum^{W}_{w}M_{i,j}^{k}[(i-\tilde{i})^{2}+(j-\tilde{j})^{2}],$
where $(\tilde{i},\tilde{j})=\arg\max_{i,j}M_{i,j}^{k}$ identifies the coordinate with the highest value in $M^{k}$, which ensuring the attribute prototype mirrors only a minimal subset of local features, thus achieving a more streamlined similarity map. This method emphasizes the precision in the prototype’s focus, leading to a denser representation of similarity. Then the $\mathcal{L}_{\text{pointwise}}$ is changed to $\mathcal{L}_{\text{pointwise}}=\sum\frac{1}{N}||a_{i}-\hat{a_{i}}||^{2}+\mathcal{L}_{\text{reg}}$. The effective representations of image locality ensure a more robust and versatile transfer of knowledge, contributing to the overall effectiveness of COMAE.
2.3. Pair-wise Attribute Consistency
To make full use of the limited yet crucial pieces of information, we conduct pair-wise attribute exploration to depict the context relationships of continuous-valued attributes. Recent studies proved that contrastive learning can enhance the discriminative capability of feature representations and enable the establishment of meaningful relationships between seen and unseen classes (Han et al., 2021; Qiu et al., 2021; Chen et al., 2018). The fundamental concept of pairwise loss revolves around learning feature representations by comparing the similarity between different pairs of samples. The pairwise loss function enhances the model’s ability to compare sample pairs across these groups, directing it to recognize that samples within the same category should exhibit higher similarity compared to those from different categories. This approach ensures a more discerning and category-aware similarity assessment within the model’s learning process. Thus we utilize contrastive learning in the pair-wise constraint. We begin with constructing the positive sample set $S^{+}_{d}(i)$ of the $i$-th image and the negative sample set $S^{-}_{d}(i)$ in the $d$-th dimension of all attributes. Specifically, given the attribute of $i$-th image, which denoted as $a_{i}$, we build the positive set of $i$-th image as follows,
(4) $S^{+}_{d}(i)=\{j\mid\lvert a_{i,d}-a_{j,d}\rvert<\epsilon,j\neq i,j=1,...,N\},$
where $\epsilon$ is a hyper-parameter and is set through multiple experiments. We randomly sample $k$ negative samples for the $i$-th image, as considering all negative samples will bring high complexity. We get the sample set $S_{d}(i)=S^{+}_{d}(i)\cup S^{-}_{d}(i)$. Afterwards, we formulate the attribute-level contrastive embedding objective as follows,
(5) $\mathcal{L}_{\text{pairwise}}=-\sum_{d}^{D}\sum_{i}^{N}\frac{1}{DN}\log\frac{\sum_{j\in S^{+}_{d}(i)}\text{exp}(\hat{a_{i}}\hat{a_{j}}/\tau)}{\sum_{t\in S_{d}(i)}\text{exp}(\hat{a_{i}}\hat{a_{t}}/\tau)},$
where $\tau>0$ is the temperature parameter. The attribute-level contrastive learning relies on attribute-wise supervision to enhance the discriminative ability in the new attribute space.
2.4. Class-wise Attribute Exploration
Class-wise attribute exploration aims to leverage image attributes to enhance the performance of classification. Compared with pointwise and pairwise explorations, it focuses on higher-level relationships of images and attributes. For high-level image representations, we adopt a global average pooling over the input image to learn a global discriminative feature $g(x)\in\mathbb{R}^{C}$, $g(x)=\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}f_{i,j}(x)$. Subsequently, a linear layer employing parameter $W\in\mathbb{R}^{C\times K}$ maps the visual feature $g(x)$ into the designated class embedding space (e.g., attribute space). Calculating the dot product between this transformed visual feature and each class embedding generates class logits. This process is complemented by applying a cross-entropy loss function, which is designed to maximize the compatibility score between the image and its corresponding attribute vector, fostering a more accurate attribute-based classification,
(6) $\mathcal{L}_{\text{ce}}=-\frac{1}{|Y^{s}|}\sum_{y_{i}\in Y^{s}}\log\frac{\text{exp}\left(g(x_{i})^{T}Wa_{i}\right)}{\text{exp}\left(g(x_{i})^{T}W\hat{a_{i}}\right)},$
The attribute embedding layer and image representations are jointly optimized, with class labels directing the overall optimization effort. Note that the attribute prototypes, which facilitate the transfer of knowledge from seen to unseen classes, are shared across different classes. This sharing mechanism enhances image representation for zero-shot learning, improving the model’s ability to generalize to new, unseen classes.
2.5. Hashing Learning
After conducting the above three consistency constraints, the image representations can contain more comprehensive attribute and class relationships. In this section, we derive a hash network based on the improved representations. Denote our hashing network as $H({\cdot})$, given an input image representation $f(x)$ from the pre-trained image encoder $f({\cdot})$. We take the Fully Connected Hash (FCH) as the fully connected layer in $f({\cdot})$ for generating binary code $b_{i}\in\mathbb{R}^{d}$, by binarization process with the use of sign function. The process is formulated as follows,
(7) $b_{i}=\text{sign}\left(\sum_{c=1}^{C}\sum_{h=1}^{H}\sum_{w=1}^{W}f(x_{i})W_{h,w,c}\right),$
where $W_{h,w,c}$ is the learnable transform matrix. It changes the last fully connected layer to a $d$-dimension representation.
To compress intra-class hashing code with strong constraints, we use the hypersphere loss (Liu et al., 2017; Wang et al., 2018) to optimize the FCH layer with large angular boundaries between classes. The spherical hashing objective offers several advantages within the context of similarity-preserving hash coding. One notable advantage lies in its ability to enforce constraints on binary codes such that they reside on a hypersphere in the Hamming space. This geometric constraint enhances the discriminative power of the learned hash codes by facilitating a more structured and separable representation of data points. Additionally, the hypersphere constraint promotes semantic similarity preservation, as points closer to the hypersphere correspond to more semantically similar items. This results in improved retrieval performance, particularly in scenarios where maintaining semantic relationships among binary codes is crucial, such as image retrieval and nearest neighbor search. Furthermore, the geometric regularity induced by the hypersphere constraint contributes to robustness against noise and variations, thereby enhancing the overall effectiveness and stability of the hashing model. It is defined as,
(8) $\mathcal{L}_{\text{hash}}$ $=-\frac{1}{N}\sum_{i}^{N}\log\left(\frac{e^{||b_{i}||\psi(\theta_{yi,i})}}{e^{||b_{i}||\psi(\theta_{yi,i})}+\sum_{j\neq y_{i}}e^{||b_{i}||\cos(\theta_{j},i)}}\right),$
$\psi(\theta_{y_{i},i})=(-1)^{k}\cos(m\theta_{y_{i},i})-2k,$
where $m\geq 1$ is an integer that controls the size of the angular margin, and $\theta_{y_{i},i}\in[\frac{k\pi}{m},\frac{(k+1)\pi}{m}],k\in[0,m-1]$. Compared with other hashing methods, it can achieve a larger separation and a smaller variation of inter-classes.
For a comprehensive exploration between the attribute learning and image hashing tasks, the overall objective is formalized as follows,
(9) $\mathcal{L}=\sum\lambda_{i}\mathcal{L}_{i},$
where $\mathcal{L}_{i}$ denotes the pointwise-, pairwise-, and classwise-losses respectively, and $\lambda_{i}$ denotes hyper-parameters. The parameters of the hashing network could not be updated by the standard back-propagation as the derivation of $\text{sign}(\cdot)$ is indifferentiable at zero. To solve this issue, the $\text{tanh}(\cdot)$ function is adopted to approximate the results of $\text{sign}(\cdot)$. In this manner, we can optimize the original indifferentiable loss by the mini-batch Stochastic Gradient Descent (SGD).
Table 1. The mAP@5000 results on three popular ZSH datasets in different code lengths.
AWA2 CUB SUN
Methods 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits
LSH 0.0106 0.0151 0.0204 0.0306 0.0055 0.0069 0.0076 0.0095 0.0598 0.0738 0.0764 0.0880
SH 0.1833 0.2729 0.2955 0.3441 0.0568 0.081 0.0886 0.1191 0.0738 0.0811 0.0837 0.0880
ITQ 0.1999 0.2821 0.2964 0.3764 0.0533 0.0765 0.0892 0.1182 0.0725 0.0899 0.0918 0.1007
IMH 0.1282 0.1536 0.1613 0.1681 0.0330 0.0361 0.0364 0.0386 0.0572 0.0668 0.0638 0.0651
PCA 0.2165 0.2530 0.2701 0.2719 0.0547 0.0598 0.0632 0.0695 0.0808 0.0868 0.0891 0.0980
HashNet 0.2086 0.2386 0.2516 0.2749 0.0528 0.0566 0.0595 0.0633 0.0746 0.0825 0.0865 0.0944
GreedyHash 0.3420 0.4169 0.4240 0.4639 0.1132 0.1707 0.1841 0.2326 0.0873 0.1259 0.1398 0.1550
JMLH 0.3607 0.4364 0.4408 0.4711 0.1078 0.1555 0.1987 0.2310 0.0860 0.1168 0.1345 0.1500
ADSH 0.3360 0.4787 0.5105 0.5454 0.0858 0.1607 0.1827 0.2424 0.0886 0.1230 0.1341 0.1731
CSQ 0.3194 0.3988 0.3773 0.4072 0.0996 0.1588 0.1712 0.2201 0.0915 0.1211 0.1375 0.1669
DPN 0.1783 0.2086 0.2378 0.2565 0.0445 0.0728 0.0772 0.1003 0.0737 0.0825 0.0894 0.1126
BiHalf 0.3440 0.4036 0.4223 0.4577 0.0794 0.1280 0.1573 0.2142 0.0077 0.0490 0.0512 0.1307
OrthoCos 0.1709 0.2302 0.2312 0.2566 0.0451 0.0660 0.0736 0.0984 0.0728 0.0809 0.0868 0.1126
CIBHash 0.2113 0.2304 0.2481 0.2618 0.0351 0.0404 0.0411 0.0453 0.0690 0.0808 0.0809 0.0883
TBH 0.0941 0.1201 0.1073 0.1730 0.0157 0.0176 0.0226 0.0252 0.0169 0.0230 0.0248 0.0287
TSK 0.2262 0.3109 0.3873 0.4151 0.0739 0.1200 0.1394 0.1112
ASZH 0.2619 0.3787 0.4032 0.4158 0.0764 0.1192 0.1294 0.1727 - - - -
SitNet 0.2344 0.2406 0.2549 0.2650 0.0880 0.1127 0.1141 0.1167 - - - -
OPZH 0.1056 0.1390 0.1618 0.1961 0.0632 0.0879 0.0962 0.1143 - - - -
AH 0.2275 0.1989 0.3154 0.3557 0.0480 0.0897 0.1089 0.1445 - - - -
COMAE 0.3819 0.4792 0.5133 0.5465 0.1136 0.1777 0.1994 0.2519 0.0923 0.1264 0.1415 0.1741
3. Experiments
In this section, we will outline the experimental setups and provide a detailed discussion of our results. Our experiments are designed to answer the following questions (RQ):
  • RQ1: Can the proposed COMAE gain superior performance in comparison to other baselines?
  • RQ2: How does the performance of COMAE evolve in response to the change in the ratio of unseen classes in the training process?
  • RQ3: What insights can be derived from an ablation study conducted on the proposed approach?
  • RQ4: How does the performance of COMAE evolve in response to the change of hyper-parameters?
3.1. Experimental Settings
(a) mAP@all in AWA2
(b) mAP@all in CUB
(c) mAP@all in SUN
(d) Decline Percentage (mAP) in AWA2
(e) Decline Percentage (mAP) in CUB
(f) Decline Percentage (mAP) in SUN
(g) AUC in AWA2
(h) AUC in CUB
(i) AUC in SUN
(j) Decline Percentage (AUC) in AWA2
(k) Decline Percentage (AUC) in CUB
(l) Decline Percentage (AUC) in SUN
Figure 3. mAP@all and AUC scores and their decline percentage with the change of unseen classes ratio in the training process.
3.1.1. Datasets
In our investigations, we meticulously analyzed three distinguished zero-shot hashing datasets referenced in (Shi et al., 2022), i.e., AWA2 (Xian et al., 2018), CUB (Wah et al., 2011), SUN (Patterson and Hays, 2012). AWA2 encompasses 37,322 images across 50 animal categories, with a configuration of 40 seen and 10 unseen classes, each described by 85 attributes [39]. The CUB dataset includes 11,788 images from 200 bird species, categorized into 150 seen and 50 unseen classes, each characterized by 312 attributes. SUN dataset, on the other hand, contains 14,340 images from 717 scene categories, with a division of 645 seen and 72 unseen classes, each depicted by 102 attributes.
3.1.2. Baselines
We select the following representative image hashing baselines as our competitors:
  • Image Shallow Hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008), SH (Weiss et al., 2008), ITQ (Gong et al., 2012), SDH (Shen et al., 2015a), IMH (Shen et al., 2015b), and PCA (Maćkiewicz and Ratajczak, 1993). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
  • Image Deep Hashing methods, leverage deep neural networks to automatically learn hierarchical and abstract representations of images. Representative works can be categorized into two paths, supervised hashing works, i.e, HashNet (Cao et al., 2017), GreedyHash (Su et al., 2018), JMLH (Shen et al., 2019), ADSH (Zhou et al., 2019), CSQ (Yuan et al., 2020), DPN (Fan et al., 2020), OrthoCos (Hoe et al., 2021), and unsupervised hashing works, i.e, BiHalf (Li and van Gemert, 2021), TBH (Qiu et al., 2021), CIBHash (Qiu et al., 2021). By employing end-to-end learning, deep hashing methods aim to optimize hash codes directly from raw pixel values, allowing for the discovery of intricate and discriminative image representations.
  • Zero-shot Image Hashing methods extend the above algorithms to handle scenarios where the model encounters classes during testing that were not seen during training. These methods often incorporate semantic information, such as image attribute descriptions, or meta-learning strategies to enhance the model’s adaptability to unseen classes. Representative works include TSK (Yang et al., 2016), ASZH, SitNet, OPZH, and AH. ASZH (Shi et al., 2022) uses category labels to learn the hash functions from the seen training data, to generate binary codes for unseen data. SitNet (Guo et al., 2017), and OPZH (Zhang et al., 2019) adopt semantic vectors to find semantic embedding space for knowledge transferability. AH (Xu et al., 2017) constructs relationships and image features for transferability.
3.1.3. Evaluation Metric
For evaluating different aspects of our study, we employed a variety of metrics. These include the mean Average Precision (mAP) and the Area Under the Precision-Recall Curve (AUC) for assessing the ranking quality of search outcomes. Specifically, we present findings for the top 5,000 samples and for the entire dataset, denoted as mAP@5000 and mAP@all, respectively. Additionally, we utilized four conventional metrics: mean Average Precision (mAP), Precision-Recall Curve (PR Curve), Precision@N Curve (P@N Curve), and Recall@N Curve (R@N Curve), with a particular focus on the ranking performance for the initial 5,000 retrieved samples.
3.1.4. Implementation Details
All comparative baselines are conducted with their default hyper-parameter configurations. For a fair comparison, their image feature extractors are substituted with the ResNet101 (Khan et al., 2018). The training epoch is set to 10 and the batch size is set to 64, and more than 10 will be overfitting, resulting in a decrease in accuracy. We take Adam as the optimizer weight decay 0.0005 and the learning rate is set to 0.0001. The $\epsilon$ in Eqn. 9 is set to 0.9, the $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ in Eqn. 4 are set to 10, 1, and 10 respectively. To ensure the statistical significance of our experimental data, the average outcomes of each technique were calculated based on ten iterations.
(a) PR Curve (AWA2)
(b) PR Curve (CUB)
(c) PR Curve (SUN)
(d) P@N Curve (AWA2)
(e) P@N Curve (CUB)
(f) P@N Curve (SUN)
(g) R@N Curve (AWA2)
(h) R@N Curve (CUB)
(i) R@N Curve (SUN)
Figure 4. The comparison of PR Curve, P@N Curve and R@N Curve in 64 bit length codes.
3.2. RQ1: Overall Performance
Table. 1 displays the results of the COMAE and all competitive baselines on datasets AWA2, CUB, and SUN with hash code lengths of 24, 48, 64, and 128. Note that the source codes of ASZH, SitNet, AH, and OPZH are unavailable, thus we use their results in paper (Shi et al., 2022). The experimental settings of Table 1 are the same as (Shi et al., 2022). In Fig. 4, we illustrate three distinct curves, i.e., PR Curve, R@N Curve, and the P@N Curve, to evaluate the performance of hash codes with a length of 64 bits on all datasets. Typically, the efficacy of a method is gauged by the PR curve approaching the upper right quadrant, with higher positions on the P@N and R@N curves indicative of superior performance. Overall, we find the following conclusions,
  • As shown in Table. 1, deep hashing models perform better than shallow hashing ones in general, which shows the superiority of neural networks. In addition, The zero-shot hashing models exhibit superior performance in comparison to both deep and shallow hashing models.
  • COMAE surpasses all existing zero-shot hashing benchmarks, in terms of all bit lengths. At the same time, the performance of COMAE remains relatively stable when the number of bits increases. Compared with AH (Xu et al., 2017), an attribute-based zero-shot hashing method, we gain better performance, which serves to validate the correctness and effectiveness of the motivation we put forth.
  • An insightful observation can be derived from the graphical representation in Fig. 4 in all datasets. COMAE attains a heightened accuracy at a comparatively lower recall rate, thereby demonstrating the superiority of our proposed method.
3.3. RQ1: Hashing Code Visualization
In this experiment, we adopted a more intuitive approach to visualizing hashing codes to evaluate the effectiveness of our proposed COMAE. Throughout the comparative experiments, we initially transformed the 64-bit hashing codes into lower-dimensional spaces, subsequently projecting these codes into a 2-D plane using the t-SNE (Van der Maaten and Hinton, 2008) tool. This methodology facilitated a clearer observation of COMAE’s performance through visual means. We present the visualization of hash code for all testing samples with unseen classes on the AWA2 dataset. Results are shown in Fig.  7. According to the visualization results, we find that a) HashNet, DPN and OrthoCos perform badly, which are unable to distinguish the unseen classes. b) Although CSQ and BiHalf gain favorable outcomes in visualization and mAP@5000 results in Table. 1, they still encounter challenges such as substantial inter-class overlap and indistinct boundaries (the classes represented by the green and other colors). c) Compared with other baselines, the proposed COMAE achieves the best performance, as the generated hash codes are more distinguished. It indicates that COMAE can effectively classify images.
(a) without $\mathcal{L}_{\text{classwise}}$
(b) with $\mathcal{L}_{\text{classwise}}$
Figure 5. Histogram distances of the intra-classes and inter-classes. The arrow annotation is the quantitative separability with the hamming distance, $\mathbb{E}[D_{inter}]-\mathbb{E}[D_{intra}]$.
3.4. RQ2: Effects of Unseen Class Ratio
To verify the robustness of our model in the context of the zero-shot scenario, we conduct experiments where the number of seen class categories underwent a progressive reduction during the training process. The results are visually depicted in Fig.  3. Specifically, (d)-(f) and (j)-(l) show the performance decrease ratio between 20%-¿40% and 20%-¿60% respectively. We conclude the following conclusions based on the outcome results,
  • All deep hashing methods manifest a discernible decline in performance concomitant with an escalation in the number of unseen classes on three datasets. Notably, this performance deterioration not only intensifies but also amplifies in magnitude, as evidenced by both the MAP@all and AUC metrics.
  • COMAE exhibits a performance superiority across various unseen class ratios, along with the most minimal decrease.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\epsilon$
Figure 6. The effects of hyper-parameters on AWA2 dataset.
(a) COMAE
(b) HashNet
(c) DPN
(d) OrthoCos
(e) CSQ
(f) BiHalf
Figure 7. t-SNE visualizations of hashing code on test unseen classes. The 10 colors denote 10 different unseen classes.
3.5. RQ3: Distance Visualization of Intra and Inter Classes
To verify whether the classwise constraint can broaden the distance of inter-class and reduce the distance of intra-class, we provide a quantitative comparison of the distance between the mean value of intra-class distances and inter-class distances, under the setting of without loss $\mathcal{L}_{\text{contrastive}}$ and with the loss. Fig. 5 summarizes the histogram of intra-class and inter-class distances on the AWA2 dataset. We adjust the frequency normalization so the total of all histogram bins equals one. The directional marker indicates the separability in Hamming distances, $\mathbb{E}[D_{inter}]-\mathbb{E}[D_{intra}]$. Despite similar distributions for inter-class distances, the distinction between the average intra-class distance (indicated by the purple dotted line) and the average inter-class distance (shown by the orange dotted line) in scenario (b) exceeds that in scenario (a). This approach visually emphasizes the effect of the classwise constraint on class separability metrics. Quantificationally, the difference in (a) is 14.38, and the difference in (b) is 15.50. The results show that our proposed classwise constraint can enlarge the distance of the inter-class and reduce the distance of the intra-class.
3.6. RQ3: Component Analysis
To provide further insight into COMAE, we conduct component analysis on AWA2 and CUB datasets with 64-bit hashing-code lengths. We configure several variants of COMAE as a) COMAE-V1 is the COMAE model without $\mathcal{L}_{\text{pointwise}}$; b) COMAE-V2 represents the COMAE without $\mathcal{L}_{\text{pairwise}}$; c) COMAE-V3 refers to the COMAE without $\mathcal{L}_{\text{classwise}}$. Results are shown in Table 2. Compared with other components, the pointwise objective plays the most significant role in AWA2, and the pairwise objective in CUB. Also, the performance of COMAE on both datasets surpasses that of variants, validating the necessity of each component.
Table 2. The mAP@all and AUC results with switching model components in different datasets.
AWA2 CUB
mAP@all AUC mAP@all AUC
COMAE-V1 0.4830 0.4781 0.2069 0.2418
COMAE-V2 0.4857 0.4791 0.1867 0.2240
COMAE-V3 0.4835 0.4782 0.2064 0.2412
COMAE 0.4883 0.4812 0.2094 0.2416
3.7. RQ4: Parameter Analysis
We analyze the impact of critical hyper-parameters, i.e., $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ that influence the weight of the auxiliary losses in Eqn. 9 and the threshold $\epsilon$ for selecting positive samples in Eqn. 4. These evaluations were performed using the AWA2 dataset with 64-bit hash codes, and the results are detailed in Fig.  6. As the increase of hyper-parameters, the performance of our model initially rises and then decreases. Hence, we choose the best hyper-parameters corresponding to the point at peak performance.
4. Conclusion
In this paper, we present COMAE for a comprehensive attribute exploration in ZSH. COMAE depicts the relationships between seeable and unseeable classes through meticulously designed explorations: point-wise, pair-wise, and class-wise constraints. Our model employs an attribute prototype network to regress attributes, facilitating the learning of local features. Then, we employ contrastive learning to capture the context of attributes, moving beyond instance-independent optimization. Finally, we introduce the class-wise constraint to cohesively enhance hashing learning, image representations, and visual attributes. Experiments demonstrate the excellent performance of COMAE.

### v5

@@
 Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios.
-While considerable success has been achieved, there still exist urgent limitations.
-Existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes.
+However, existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes.
 Also, the continuous-value attributes are not fully harnessed.
-In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints.
+In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints.
 By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes.
@@
 Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively.
-Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes.
-Our codes are available at https://github.com/itsnotacie/TOMM2024-COMAE.
-Abstract.
-Zero-shot hashing (ZSH) has shown excellent success owing to its efficiency and generalization in large-scale retrieval scenarios.
-While considerable success has been achieved, there still exist urgent limitations.
-Existing works ignore the locality relationships of representations and attributes, which have effective transferability between seeable classes and unseeable classes.
-Also, the continuous-value attributes are not fully harnessed.
-In response, we conduct a COMprehensive Attribute Exploration for ZSH, named COMAE, which depicts the relationships from seen classes to unseen ones through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints.
-By regressing attributes from the proposed attribute prototype network, COMAE learns the local features that are relevant to the visual attributes.
-Then COMAE utilizes contrastive learning to comprehensively depict the context of attributes, rather than instance-independent optimization.
-Finally, the class-wise constraint is designed to cohesively learn the hash code, image representation, and visual attributes more effectively.
-Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes.
-Our source codes are available at https://github.com/YihangZhou0424/NN2024-COMAE.
-Prototype Learning, Local Representation, Zero-shot Hashing, Image Zero-shot Hashing, Image Hashing, Image Retrieval.
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXjournal: TOMMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Computing methodologies Visual content-based indexing and retrievalccs: Information systems Image search
-Image hashing techniques have garnered substantial interest due to their remarkable efficacy in multimedia retrieval and content-based searches for images and documents (Lew et al., 2006; Wang et al., 2017; Ju et al., 2024).
-Particularly, hashing techniques convert high-dimensional and complicated data into binary codes, which enables efficient searches in the binary Hamming space, thereby contributing to a smaller memory footprint and increased search efficiency (Shi et al., 2022; Li and van Gemert, 2021; Ju et al., 2023).
-The capability to depict intricate data in a compact binary format not only accelerates retrieval processes but also markedly improves storage efficiency.
-Consequently, image hashing emerges as a highly valuable asset in situations that necessitate swift and resource-efficient information retrieval.
-With the popularity of content and social apps, the growth of Internet images leads to sort of new concepts, presenting a challenge for traditional hashing algorithms that struggle to adapt to unseen data in previously unobserved categories (Venkatesan et al., 2000; Wang et al., 2017; Long et al., 2021).
+Furthermore, a theoretical analysis is provided to show the effectiveness.
+Experimental results demonstrate that COMAE outperforms state-of-the-art hashing models, especially in scenarios with a larger number of unseen label classes.
+Our codes are available at https://github.com/itsnotacie/ICMR2025-COMAE.
+Deep Hash, Zero-shot Hashing, Attribute Learning, Prototype Learning, Image Retrieval, Image Hashing
+journalyear: 2025copyright: ccconference: Proceedings of the 2025 International Conference on Multimedia Retrieval; June 30-July 3, 2025; Chicago, IL, USAbooktitle: Proceedings of the 2025 International Conference on Multimedia Retrieval (ICMR ’25), June 30-July 3, 2025, Chicago, IL, USAdoi: 10.1145/3731715.3733294isbn: 979-8-4007-1877-9/2025/06ccs: Information Systems Similarity Measuresccs: Information Systems Information retrievalccs: Information systems Image searchccs: Computing methodologies Visual content-based indexing and retrieval
+1. Introduction
+Image hashing techniques have garnered substantial interest due to their remarkable efficacy in multimedia retrieval and content-based searches for images and documents (Wang et al., 2017; Dong et al., 2024a; Arik and Pfister, 2019; Luo et al., 2025; Long et al., 2021; Li et al., 2025c, d; Zeng et al., 2025).
+Particularly, hashing techniques convert high-dimensional and complicated data into binary codes, which enables efficient searches in the binary Hamming space, thereby contributing to a smaller memory footprint and increased search efficiency (Shi et al., 2022; Li and van Gemert, 2021; Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019; Long et al., 2020).
+With the popularity of content and social apps, the growth of Internet images leads to sort of new concepts, presenting a challenge for traditional hashing algorithms that struggle to adapt to unseen data in previously unobserved categories (Venkatesan et al., 2000; Wang et al., 2017).
 Recent researchers paid attention to the Zero-Shot Hashing (ZSH) issue.
-Visual attributes, pervasive across prevalent image datasets, serve to delineate discriminative visual properties intrinsic to objects, transcending specific classes and establishing a foundation for addressing the intricacies of previously unseen data (Khan et al., 2018; Chen et al., 2022).
+Visual attributes, pervasive across prevalent image datasets, delineate discriminative visual properties intrinsic to objects, transcending specific classes and establishing a foundation for addressing the intricacies of previously unseen data (Khan et al., 2018; Chen et al., 2022; Ma et al., 2024).
 Notably, these attributes offer a robust representation for recognizing out-of-distribution samples, making them valuable for capturing similarities in the seen and unseen concepts in Zero-Shot Learning (ZSL) tasks (Arik and Pfister, 2019; Xu et al., 2017).
-Fig.
-1 shows an example of the class-level attributes.
 Nevertheless, the exploration of visual attributes within the realm of ZSH tasks has been relatively meager.
 AH (Xu et al., 2017) adopts a straightforward approach by constructing a multi-task architecture to capture the intricate relationships between visual attributes and category classes, thereby facilitating the transfer of knowledge.
-While considerable success has been achieved in utilizing visual attributes to establish connections among visual attributes, image features, class labels, and binary hash codes, there still exists urgent limitations (Li and van Gemert, 2021; Xu et al., 2017).
-Firstly, image locality representations are proven to be critical to zero-shot learning tasks (Li et al., 2021), as the locality exhibits more effective transferability between seen classes and unseen classes.
+While considerable success has been achieved in utilizing visual attributes to establish connections among visual attributes, image features, class labels, and binary hash codes, there still exist urgent limitations (Li and van Gemert, 2021; Xu et al., 2017).
+Firstly, image locality representations are proven to be critical to zero-shot learning tasks (Li et al., 2021a), as the locality exhibits more effective transferability between seen classes and unseen classes.
 However, existing works ignore the locality representations in deep hashing tasks, which decreases the upper bound of the transferability in hashing algorithms.
@@
 Thirdly, existing works generally employ a multi-task learning framework to independently utilize attribute labels and class labels.
-Yet their actual relationships are more complicated than such simple presumes.
+Yet their actual relationships are more complicated than such simple presumptions.
 To be specific, two images may belong to the same class but pose disparate attributes.
-For example, two images with a “bird” class may have “black leg” and “yellow leg”.
-Similarly, instances with different classes may also exhibit sharing analogous attributes.
+For example, as shown in Fig.
+1, two images with a “bird” class may have “black leg” and “brown leg”.
+Similarly, instances with different classes may also exhibit analogous attributes.
 For example, two images with “bird” and “wapiti” classes may have the same “black leg”.
@@
 Extensive experiments on the ZSH benchmark datasets show that COMAE outperforms the state-of-the-art methods, especially in scenarios with a larger number of unseen label classes.
-It serves to validate the effectiveness of comprehensive attribute explorations.
-Figure 1. An example of the class-level attributes.
 Our contributions are summarized as follows:
-  • We conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE, it depicts the relationships between seen classes and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise and class-wise consistency constraints.
-  • We design an attribute prototype network, a contrastive learning unit, and a joint-optimal label utilization to implement the above three components.
+  • We conduct a COMprehensive Attribute Exploration for zero-shot hashing, named COMAE, which depicts the relationships between seen classes and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints.
+  • We design an attribute prototype network, a contrastive learning unit, and a joint-optimal label utilization to implement the above three components, with theoretical analysis to show its effectiveness.
   • Experiments on three popular zero-shot hashing datasets show that COMAE significantly outperforms the state-of-the-art methods, especially in scenarios with a larger number of unseen label classes.
-This paper is generally organized as follows.
-We begin with briefly reviewing related works in Section 2.
-Then the proposed COMAE is technically introduced in Section 3.
-We provide extensive experiment results and discussions in Section 4.
-Finally, we conclude this work in Section 5.
+Figure 1.
+The architecture of the proposed COMAE.
+It consists of three modules: a) Point-wise objective aims at improving image locality and attribute representations; b) Pair-wise loss is proposed to learn the representations from individual learning to context-based learning; and c) Class-wise constraint is designed to capture relationships of attributes and class labels.
 2. Related Work
 2.1. Deep Hashing
-Image hashing is a popularly used method for approximating nearest neighbor search in large-scale retrieval scenarios (Venkatesan et al., 2000).
+Image hashing is a popularly used method for approximating nearest neighbor search in large-scale retrieval scenarios (Venkatesan et al., 2000; Dong et al., 2024b; Qin et al., 2024; Chen et al., 2025a).
 Shallow image hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations.
@@
 Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
-Over the past decades, deep learning models have shown excellent progress in the field of Computer Vision (Khan et al., 2018; Pourpanah et al., 2022) and Natural Language Processing (Song et al., 2020; Long et al., 2021), many scholars pay attention to leveraging deep neural networks to learn the abstract representations.
+Over the past decades, deep learning models have shown excellent progress in the field of computer vision (Zhu et al., 2024a, b; Dong et al., 2024a; Chu et al., 2025; Qin et al., 2023), language processing (Wu et al., 2024; Guo et al., 2025; Cai et al., 2023; Xiao et al., 2025; Lu et al., 2024), and multimodal tasks (Lin et al., 2024; Li et al., 2021b; Xiao et al., 2021; Li et al., 2024, 2025b).
 Deep hashing methods can be achieved through three paths.
 The first path utilizes limited information, which has a lower upper bound of the capability.
-ASZH (Shi et al., 2022) is a typical work in the first path, that learns hash functions based on the seeable category labels, for generating binary codes for unseeable data.
-The second path (Zhang et al., 2019; Guo et al., 2017) generates semantic vectors based on the visual attributes to find a semantic embedding space, then transfers the common knowledge to unseeable classes.
-TSK, SitNet, and OPZH (Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019) are representative semantic-based works, they conduct a multi-task framework in which the class labels and attribute vectors are co-trained through the hashing network, and the labels are projected into a semantic embedding space.
+ASZH (Shi et al., 2022; Qin et al., 2022) is a typical work in the first path, which learns hash functions based on the visible category labels, for generating binary codes for unseeable data.
+The second path (Zhang et al., 2019; Guo et al., 2017) generates semantic vectors based on the visual attributes to find a semantic embedding space, then transfers the common knowledge to unseen classes.
+TSK, SitNet, OPZH, and PIXEL (Yang et al., 2016; Guo et al., 2017; Zhang et al., 2019; Dong et al., 2024a) are representative semantic-based works, they conduct a multi-task framework in which the class labels and attribute vectors are co-trained through the hashing network, and the labels are projected into a semantic embedding space.
+However, PIXEL introduced BERT, which resulted in longer processing times.
 However, there exists an inherent gap between the semantic and visual space, and the alignment of spaces is tough.
@@
 2.2. Zero-shot Learning
-Zero-shot learning (ZSL) (Farhadi et al., 2009; Xu et al., 2020; Zheng and Gupta, 2022; Long et al., 2020) is designed to extend object recognition across known and unknown classes via a unified embedding space, where both categories are described by their semantic visual identifiers.
-Early ZSL methods (Farhadi et al., 2009; Pourpanah et al., 2022) harvest global visual features from networks, whether pre-trained or capable of end-to-end training.
+Zero-shot learning (ZSL) (Farhadi et al., 2009; Xu et al., 2020; Zheng and Gupta, 2022) is designed to extend object recognition across known and unknown classes via a unified embedding space, where both categories are described by their semantic visual identifiers.
+Early ZSL methods (Farhadi et al., 2009; Pourpanah et al., 2022; Chen et al., 2025b) harvest global visual features from networks, whether pre-trained or capable of end-to-end training.
 Typically, end-to-end frameworks often surpass pre-trained ones by refining visual attributes, thus mitigating biases between the datasets used for training and those for ZSL applications.
-Nonetheless,such strategies typically produce less than ideal outcomes (Xu et al., 2020), as they struggle to discern the subtle variances between seen and unseen categories.
+Nonetheless,such strategies typically produce less than ideal outcomes (Xu et al., 2020; Li et al., 2025a), as they struggle to discern the subtle variances between seen and unseen categories.
 Recent innovations in ZSL predominantly concentrate on methods based on attention (Zheng and Gupta, 2022; Sylvain et al., 2019; Xu et al., 2020), employing visual attributes to identify distinct feature regions, for instance, specific parts of a bird.
@@
 Inspired by these insights into visual attributes and the delicate distinctions between recognized and unrecognized classes, our methodology adopts the prototype network to intricately map the relationships between visual attributes and hash codes (Xu et al., 2020).
-Figure 2.
-The architecture of the proposed COMAE.
-It consists of three modules: a) Point-wise objective aims at improving image locality and attribute representations; b) Pair-wise loss is proposed to learn the representations from individual learning to context-based learning; and c) Class-wise constraint is designed to capture relationships of attributes and class labels.
-1.3. Contrastive Learning
-Contrastive learning aims at honing representations by juxtaposing similar instance pairs (positive pairs) against dissimilar ones (negative pairs) (Hadsell et al., 2006; Han et al., 2021), enhancing the model’s ability to differentiate between closely related and distinct instances effectively.
-Many works demonstrate that contrastive learning is able to alleviate the problems of numerous categories and imbalanced class distributions.
-In addition, it can amplify the subtle distinctions among similar classes, which assists in capturing intricate details that are often overlooked by basic models.
-(Hadsell et al., 2006) is the first attempt work to improve learned representations by contrasting positive and negative pairs.
-SimCLR (Chen et al., 2020) utilizes a self-supervised learning architecture that substitutes the conventional memory bank with batch-specific elements, significantly enhancing its performance on the ImageNet dataset.
-As hash code is a form of representation, recent researches (Yu et al., 2022; Song et al., 2023; Qiu et al., 2021) take attempts to utilize contrastive learning in deep hashing.
-(Yu et al., 2022) proposes weighted contrastive hashing to explore the fine-grained semantic relations among images.
-(Song et al., 2023) propose to However, random sampling could result in augmented images with different semantics, deteriorating the performance of contrastive learning.
-2. Method
+3. Model: COMAE
 The architecture of the proposed COMAE is shown in Fig.
-2.
+1.
 Our proposed COMAE depicts the relationships between seen and unseen classes through three meticulously designed explorations, i.e., point-wise, pair-wise, and class-wise consistency constraints.
 To improve the image locality and attribute representations, we propose a point-wise objective.
-To learn the representations from individual learning to context-based learning, we utilize contrastive learning to build pair-wise loss.
+To learn the representations from individual learning to context-based learning, we utilize contrastive learning to build a pair-wise loss.
 For capturing relationships between attributes and class labels, a class-wise constraint is designed.
 3.1. Problem Definition
-Let $X=\{x_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times H\times W\times C}$ be $N$ images, where $H,W,C$ denote the image height, width, and channel.
-The matrix $Y=\{y_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C}$ is the label matrix, where $C$ denotes the number of classes.
-The classes are divided into the seeable class $Y^{s}$ and the unseeable class $Y^{u}$, where $|Y^{s}|\cup|Y^{u}|=|C|,|Y^{s}|\cap|Y^{u}|=\emptyset$.
+Let $X=\{x_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times H\times W\times C}$ be $N$ images, where $H,W,C$ denote the image height, width and channel.
+The matrix $Y=\{y_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C}$ denotes the labels, where $C$ denotes the number of classes.
+The labels are divided into the seeable class $Y^{s}$ and the unseeable class $Y^{u}$, where $|Y^{s}|\cup|Y^{u}|=|C|,|Y^{s}|\cap|Y^{u}|=\emptyset$.
 Let $A=\{a_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times K}$ denote the class-level attribute vector in $K$ dimensions.
-By using training samples which contain labeled images and attributes from seeable classes, i.e., $S=\{x,y,a|x\in X,y\in Y^{s},a\in A\}$, we aim to learn a hashing model $\mathcal{H}\colon X\rightarrow b_{i}\in\{-1,1\}^{l}$, where $b_{i}$ represents the learned $l$-bit binary hash codes.
-In the testing process, attribute vectors of unseen classes are also known.
+By using training samples which contain labeled images and attributes from seeable classes, i.e., $S=\{x,y,a|x\in X,y\in Y^{s},a\in A\}$, we aim to learn a hashing model $\mathcal{H}\colon X\rightarrow b_{i}\in\{-1,1\}^{l}$, where $b_{i}$ denotes learned $l$-bit binary hash codes.
+For the testing process, attribute vectors of unseen classes are also known.
 In this paper, zero-shot hashing focuses on predicting the hashing label of images from unseen classes, i.e., $X\rightarrow Y^{u}$.
 3.2. Point-wise Attribute Consistency
-Image locality representations are proven to be practical in transferring knowledge from seen to unseen classes (Xu et al., 2020; Li et al., 2021).
-For more subtly depicted image and attribute characterizations, we first conduct point-wise consistency constraint, which takes the attribute regression loss as the supervision.
+Image locality representations are proven to be practical in transferring knowledge from seen to unseen classes (Xu et al., 2020; Li et al., 2021a).
+For more subtly depicted image and attribute characterizations, we first conduct a point-wise consistency constraint, which takes the attribute regression loss as the supervision.
 Prototype learning has proven to be robust in dealing with open-set recognition and out-of-distribution samples (Xu et al., 2020; Tan et al., 2022) in recent years.
@@
 As shown in Fig.
-2, $p_{1}$ and $p_{2}$ correspond to the prototypes for “red beak” and “yellow leg” respectively.
-For each attribute, we produce a similarity map $M^{k}\in\mathbb{R}^{H\times W}$, where each element is computed by a non-linear network $\Theta(\cdot)$ between for $p_{k}$ and each local feature.
+1, $p_{1}$ and $p_{2}$ correspond to the prototypes for “orange crown” and “yellow skin” respectively.
+For each attribute, we produce a similarity map $M^{k}\in\mathbb{R}^{H\times W}$, where each element is computed by a non-linear network $\Theta(\cdot)$ between $p_{k}$ and each local feature.
 To associate each visual attribute with its closest local feature, we gain the predicted $k$-th attribute $\hat{a_{k}}$ by taking the maximum value in the similarity map $M^{k}$,
@@
 (2) $\mathcal{L}_{\text{pointwise}}=\sum_{i}^{N}\frac{1}{N}||a_{i}-\hat{a_{i}}||^{2}.$
-Furthermore, we refine the similarity map to intensively focus on its peak region.
-To achieve this, we employ a compactness regularizer (Zheng et al., 2017) on each similarity map $M^{k}$, ensuring a concentrated similarity assessment.
-The equation is defined as,
-(3) $\mathcal{L}_{\text{reg}}=\sum^{K}_{k}\sum^{H}_{h}\sum^{W}_{w}M_{i,j}^{k}[(i-\tilde{i})^{2}+(j-\tilde{j})^{2}],$
-where $(\tilde{i},\tilde{j})=\arg\max_{i,j}M_{i,j}^{k}$ identifies the coordinate with the highest value in $M^{k}$, which ensuring the attribute prototype mirrors only a minimal subset of local features, thus achieving a more streamlined similarity map.
-This method emphasizes the precision in the prototype’s focus, leading to a denser representation of similarity.
-Then the $\mathcal{L}_{\text{pointwise}}$ is changed to $\mathcal{L}_{\text{pointwise}}=\sum\frac{1}{N}||a_{i}-\hat{a_{i}}||^{2}+\mathcal{L}_{\text{reg}}$.
-The effective representations of image locality ensure a more robust and versatile transfer of knowledge, contributing to the overall effectiveness of COMAE.
 3.3. Pair-wise Attribute Consistency
@@
 (3) $S^{+}_{d}(i)=\{j\mid\lvert a_{i,d}-a_{j,d}\rvert<\epsilon,j\neq i,j=1,...,N\},$
-where $\epsilon$ is a hyper-parameter and is set through multiple experiments.
+where $\epsilon$ is a hyperparameter and is set through multiple experiments.
 We randomly sample $k$ negative samples for the $i$-th image, as considering all negative samples will bring high complexity.
@@
 where $\tau>0$ is the temperature parameter. The attribute-level contrastive learning relies on attribute-wise supervision to enhance the discriminative ability in the new attribute space.
+Algorithm 1 Algorithm of COMAE
+1:Training Images $X=\{x_{i}\}_{i=1}^{N}$, class labels $Y=\{y_{i}\}_{i=1}^{N}$, and image attribute vectors $A=\{a_{i}\}_{i=1}^{N}$.
+2:The parameters $\Theta$
+3:for $i=0,...,N$ do
+4:     Generate image initial embeddings $f(x_{i})$ based on the image encoder, i.e., ResNet101;
+5:end for
+6:Generate the predicted prototypes $\{M^{k}\}_{k=1}^{K}$ for each attribute.
+7:for $i=0,...,N$ do
+8:     Construct the positive sample set $S^{+}_{d}(i)$ of the $i$-th image and negative sample set $S^{-}_{d}(i)$, as Eqn. (3).
+9:end for
+10:while not converge do :
+11:     Compute the pointwise loss, as Eqn. (2).
+12:     Compute the pairwise loss, as Eqn. (4).
+13:     Compute the classwise loss, as Eqn. (5).
+14:     for $i=0,...,N$ do
+15:         Generate binary hashing codes $B=\{b_{i}\}_{i=1}^{N}$ based on image representations $f(x_{i})$ through Eqn. (6).
+16:     end for
+17:     Compute the joint loss through Eqn. (7)
+18:     Compute the backpropagation error.
+19:     Update the parameters $\Theta$ and image hidden representations $f(x_{i})$.
+20:end while
+21:return Image hashing codes $B=\{b_{i}\}_{i=1}^{N},b_{i}\in\{-1,1\}^{l}$.
 3.4. Class-wise Attribute Exploration
@@
 (6) $b_{i}=\text{sign}\left(\sum_{c=1}^{C}\sum_{h=1}^{H}\sum_{w=1}^{W}f(x_{i})W_{h,w,c}\right),$
-where $W_{h,w,c}$ is the learnable transform matrix. It changes the last fully connected layer to a $d$-dimension representation.
-To compress intra-class hashing code with strong constraints, we use the hypersphere loss (Liu et al., 2017; Wang et al., 2018) to optimize the FCH layer with large angular boundaries between classes.
-The spherical hashing objective offers several advantages within the context of similarity-preserving hash coding.
-One notable advantage lies in its ability to enforce constraints on binary codes such that they reside on a hypersphere in the Hamming space.
-This geometric constraint enhances the discriminative power of the learned hash codes by facilitating a more structured and separable representation of data points.
-Additionally, the hypersphere constraint promotes semantic similarity preservation, as points closer to the hypersphere correspond to more semantically similar items.
-This results in improved retrieval performance, particularly in scenarios where maintaining semantic relationships among binary codes is crucial, such as image retrieval and nearest neighbor search.
-Furthermore, the geometric regularity induced by the hypersphere constraint contributes to robustness against noise and variations, thereby enhancing the overall effectiveness and stability of the hashing model.
-It is defined as,
-(8) $\mathcal{L}_{\text{hash}}$ $=-\frac{1}{N}\sum_{i}^{N}\log\left(\frac{e^{||b_{i}||\psi(\theta_{yi,i})}}{e^{||b_{i}||\psi(\theta_{yi,i})}+\sum_{j\neq y_{i}}e^{||b_{i}||\cos(\theta_{j},i)}}\right),$
-$\psi(\theta_{y_{i},i})=(-1)^{k}\cos(m\theta_{y_{i},i})-2k,$
-where $m\geq 1$ is an integer that controls the size of the angular margin, and $\theta_{y_{i},i}\in[\frac{k\pi}{m},\frac{(k+1)\pi}{m}],k\in[0,m-1]$.
+where $W_{h,w,c}$ is the learnable transform matrix.
+It changes the last fully connected layer to a $d$-dimension representation.
+To compress intra-class hashing code with strong constraints, we use the hypersphere loss (Liu et al., 2017; Wang et al., 2018) to optimize the FCH layer with large angular boundaries between classes as $\mathcal{L}_{\text{hash}}$.
 Compared with other hashing methods, it can achieve a larger separation and a smaller variation of inter-classes.
-For a comprehensive exploration between the attribute learning and image hashing tasks, the overall objective is formalized as follows,
+For a comprehensive exploration between the attribute learning and hashing tasks, the overall objective is formalized as follows,
 (7) $\mathcal{L}=\sum\lambda_{i}\mathcal{L}_{i},$
-where $\mathcal{L}_{i}$ denotes the pointwise-, pairwise-, and classwise-losses respectively, and $\lambda_{i}$ denotes hyper-parameters.
+where $\mathcal{L}_{i}$ denotes the pointwise-, pairwise-, classwise-, and hashing losses respectively, and $\lambda_{i}$ denotes hyper-parameters.
 The parameters of the hashing network could not be updated by the standard back-propagation as the derivation of $\text{sign}(\cdot)$ is indifferentiable at zero.
-To solve this issue, the $\text{tanh}(\cdot)$ function is adopted to approximate the results of $\text{sign}(\cdot)$.
-In this manner, we can optimize the original indifferentiable loss by the mini-batch Stochastic Gradient Descent (SGD).
+Thus $\text{tanh}(\cdot)$ is adopted to approximate the results of $\text{sign}(\cdot)$.
+In this manner, we can optimize the original non-differentiable loss by the mini-batch Stochastic Gradient Descent (SGD).
+The pseudocode of COMAE is shown in the Algorithm.
+1.
+3.6. Theoretical Analysis
+Here we provide a theoretical analysis for our proposed COMAE.
+Theorem 1.
+[…]
+[…]
+[…]
+Theorem 2.
+[…]
+[…]
 Table 1. The mAP@5000 results on three popular ZSH datasets in different code lengths.
 AWA2 CUB SUN
-Methods 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits
-LSH 0.0106 0.0151 0.0204 0.0306 0.0055 0.0069 0.0076 0.0095 0.0598 0.0738 0.0764 0.0880
-SH 0.1833 0.2729 0.2955 0.3441 0.0568 0.081 0.0886 0.1191 0.0738 0.0811 0.0837 0.0880
-ITQ 0.1999 0.2821 0.2964 0.3764 0.0533 0.0765 0.0892 0.1182 0.0725 0.0899 0.0918 0.1007
-IMH 0.1282 0.1536 0.1613 0.1681 0.0330 0.0361 0.0364 0.0386 0.0572 0.0668 0.0638 0.0651
-PCA 0.2165 0.2530 0.2701 0.2719 0.0547 0.0598 0.0632 0.0695 0.0808 0.0868 0.0891 0.0980
-HashNet 0.2086 0.2386 0.2516 0.2749 0.0528 0.0566 0.0595 0.0633 0.0746 0.0825 0.0865 0.0944
-GreedyHash 0.3420 0.4169 0.4240 0.4639 0.1132 0.1707 0.1841 0.2326 0.0873 0.1259 0.1398 0.1550
-JMLH 0.3607 0.4364 0.4408 0.4711 0.1078 0.1555 0.1987 0.2310 0.0860 0.1168 0.1345 0.1500
-ADSH 0.3360 0.4787 0.5105 0.5454 0.0858 0.1607 0.1827 0.2424 0.0886 0.1230 0.1341 0.1731
-CSQ 0.3194 0.3988 0.3773 0.4072 0.0996 0.1588 0.1712 0.2201 0.0915 0.1211 0.1375 0.1669
-DPN 0.1783 0.2086 0.2378 0.2565 0.0445 0.0728 0.0772 0.1003 0.0737 0.0825 0.0894 0.1126
-BiHalf 0.3440 0.4036 0.4223 0.4577 0.0794 0.1280 0.1573 0.2142 0.0077 0.0490 0.0512 0.1307
-OrthoCos 0.1709 0.2302 0.2312 0.2566 0.0451 0.0660 0.0736 0.0984 0.0728 0.0809 0.0868 0.1126
-CIBHash 0.2113 0.2304 0.2481 0.2618 0.0351 0.0404 0.0411 0.0453 0.0690 0.0808 0.0809 0.0883
-TBH 0.0941 0.1201 0.1073 0.1730 0.0157 0.0176 0.0226 0.0252 0.0169 0.0230 0.0248 0.0287
-TSK 0.2262 0.3109 0.3873 0.4151 0.0739 0.1200 0.1394 0.1112
-ASZH 0.2619 0.3787 0.4032 0.4158 0.0764 0.1192 0.1294 0.1727 - - - -
-SitNet 0.2344 0.2406 0.2549 0.2650 0.0880 0.1127 0.1141 0.1167 - - - -
-OPZH 0.1056 0.1390 0.1618 0.1961 0.0632 0.0879 0.0962 0.1143 - - - -
-AH 0.2275 0.1989 0.3154 0.3557 0.0480 0.0897 0.1089 0.1445 - - - -
-COMAE 0.3819 0.4792 0.5133 0.5465 0.1136 0.1777 0.1994 0.2519 0.0923 0.1264 0.1415 0.1741
+Methods Reference 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits
+LSH CACM08 0.0106 0.0151 0.0204 0.0306 0.0055 0.0069 0.0076 0.0095 0.0598 0.0738 0.0764 0.0880
+SH NeurIPS08 0.1833 0.2729 0.2955 0.3441 0.0568 0.081 0.0886 0.1191 0.0738 0.0811 0.0837 0.0880
+ITQ TPAMI12 0.1999 0.2821 0.2964 0.3764 0.0533 0.0765 0.0892 0.1182 0.0725 0.0899 0.0918 0.1007
+IMH TIP15 0.1282 0.1536 0.1613 0.1681 0.0330 0.0361 0.0364 0.0386 0.0572 0.0668 0.0638 0.0651
+PCA CG93 0.2165 0.2530 0.2701 0.2719 0.0547 0.0598 0.0632 0.0695 0.0808 0.0868 0.0891 0.0980
+HashNet ICCV17 0.2086 0.2386 0.2516 0.2749 0.0528 0.0566 0.0595 0.0633 0.0746 0.0825 0.0865 0.0944
+GreedyHash NeurIPS18 0.3420 0.4169 0.4240 0.4639 0.1132 0.1707 0.1841 0.2326 0.0873 0.1259 0.1398 0.1550
+JMLH ICCV19 0.3607 0.4364 0.4408 0.4711 0.1078 0.1555 0.1987 0.2310 0.0860 0.1168 0.1345 0.1500
+ADSH Access19 0.3360 0.4787 0.5105 0.5454 0.0858 0.1607 0.1827 0.2424 0.0886 0.1230 0.1341 0.1731
+CSQ CVPR20 0.3194 0.3988 0.3773 0.4072 0.0996 0.1588 0.1712 0.2201 0.0915 0.1211 0.1375 0.1669
+DPN IJCAI20 0.1783 0.2086 0.2378 0.2565 0.0445 0.0728 0.0772 0.1003 0.0737 0.0825 0.0894 0.1126
+BiHalf AAAI21 0.3440 0.4036 0.4223 0.4577 0.0794 0.1280 0.1573 0.2142 0.0077 0.0490 0.0512 0.1307
+OrthoCos NeurIPS21 0.1709 0.2302 0.2312 0.2566 0.0451 0.0660 0.0736 0.0984 0.0728 0.0809 0.0868 0.1126
+CIBHash IJCAI21 0.2113 0.2304 0.2481 0.2618 0.0351 0.0404 0.0411 0.0453 0.0690 0.0808 0.0809 0.0883
+TBH CVPR20 0.0941 0.1201 0.1073 0.1730 0.0157 0.0176 0.0226 0.0252 0.0169 0.0230 0.0248 0.0287
+TSK MM16 0.2262 0.3109 0.3873 0.4151 0.0739 0.1200 0.1394 0.1112
+ASZH TKDE22 0.2619 0.3787 0.4032 0.4158 0.0764 0.1192 0.1294 0.1727 - - - -
+SitNet IJCAI17 0.2344 0.2406 0.2549 0.2650 0.0880 0.1127 0.1141 0.1167 - - - -
+OPZH PRL19 0.1056 0.1390 0.1618 0.1961 0.0632 0.0879 0.0962 0.1143 - - - -
+AH ICME17 0.2275 0.1989 0.3154 0.3557 0.0480 0.0897 0.1089 0.1445 - - - -
+PIXEL CIKM24 0.3981 0.4792 0.5133 0.5465 0.1136 0.1777 0.1994 0.2519 0.0923 0.1264 0.1415 0.1741
+COMAE Ours 0.3843 0.5073 0.5149 0.5801 0.1142 0.1803 0.2012 0.2532 0.0925 0.1278 0.1465 0.1764
 4. Experiments
 In this section, we will outline the experimental setups and provide a detailed discussion of our results. Our experiments are designed to answer the following questions (RQ):
-  • RQ1: Can the proposed COMAE gain superior performance in comparison to other baselines?
-  • RQ2: How does the performance of COMAE evolve in response to the change in the ratio of unseen classes in the training process?
-  • RQ3: What insights can be derived from an ablation study conducted on the proposed approach?
-  • RQ4: How does the performance of COMAE evolve in response to the change of hyper-parameters?
+  • RQ1: Can COMAE gain superior performance in comparison to other baselines?
+  • RQ2: How does COMAE perform in response to the change in the ratio of unseen classes in the training process?
+  • RQ3: How does COMAE perform in case studies, where its effectiveness is evaluated in a more intuitive way?
+  • RQ4: Is COMAE a running efficient hashing model?
+  • RQ5: How does COMAE perform in model analysis, including the evaluation of its component change?
 4.1. Experimental Settings
+4.1.1. Datasets
+In this paper, we analyze three distinguished zero-shot hashing datasets referenced in (Shi et al., 2022), i.e., AWA2 (Xian et al., 2018), CUB (Wah et al., 2011), SUN (Patterson and Hays, 2012).
+AWA2 encompasses 37,322 images across 50 animal categories, with a configuration of 40 seen and 10 unseen classes, each described by 85 attributes [39].
+The CUB dataset includes 11,788 images from 200 bird species, categorized into 150 seen and 50 unseen classes, each characterized by 312 attributes.
+SUN dataset, on the other hand, contains 14,340 images from 717 scene categories, with a division of 645 seen and 72 unseen classes, each depicted by 102 attributes.
+4.1.2. Baselines
+We select the following representative image hashing baselines as our competitors:
+(1) Image Shallow Hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations.
+Typical works include LSH (Andoni and Indyk, 2008), SH (Weiss et al., 2008), ITQ (Gong et al., 2012), IMH (Shen et al., 2015), and PCA (Maćkiewicz and Ratajczak, 1993).
+Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
+(2) Image Deep Hashing methods, leverage deep neural networks to automatically learn hierarchical and abstract representations of images.
+Representative works can be categorized into two paths, supervised hashing works, i.e, HashNet (Cao et al., 2017), GreedyHash (Su et al., 2018), JMLH (Shen et al., 2019), ADSH (Zhou et al., 2019), CSQ (Yuan et al., 2020), DPN (Fan et al., 2020), OrthoCos (Hoe et al., 2021), and unsupervised hashing works, i.e, BiHalf (Li and van Gemert, 2021), TBH (Shen et al., 2020), CIBHash (Qiu et al., 2021).
+By employing end-to-end learning, deep hashing methods aim to optimize hash codes directly from raw pixel values, allowing for the discovery of intricate and discriminative image representations.
+(3) Zero-shot Image Hashing methods extend the above algorithms to handle scenarios where the model encounters classes during testing that were not seen during training.
+These methods often incorporate semantic information, such as image attribute descriptions, or meta-learning strategies to enhance the model’s adaptability to unseen classes.
+Representative works include TSK (Yang et al., 2016), ASZH, SitNet, OPZH, and AH, PIXEL(Dong et al., 2024a).
+ASZH (Shi et al., 2022) uses category labels to learn the hash functions from the seen training data, to generate binary codes for unseen data.
+SitNet (Guo et al., 2017), and OPZH (Zhang et al., 2019) adopt semantic vectors to find a semantic embedding space for knowledge transferability.
+AH (Xu et al., 2017) constructs relationships and image features for transferability.
+PIXEL(Dong et al., 2024a) improves accuracy by incorporating BERT as the text encoder.
+4.1.3. Evaluation Metric
+To evaluate different aspects of our study, we employed a variety of metrics.
+These include the mean Average Precision (mAP) and the Area Under the Precision-Recall Curve (AUC) for assessing the ranking quality of search outcomes.
+Specifically, we present findings for the top 5,000 samples and for the entire dataset, denoted as mAP@5000 and mAP@all, respectively.
+Additionally, we utilized four conventional metrics: mean Average Precision (mAP), Precision-Recall Curve (PR Curve), Precision@N Curve (P@N Curve), and Recall@N Curve (R@N Curve), with a particular focus on the ranking performance for the initial 5,000 retrieved samples.
+4.1.4. Implementation Details
+All comparative baselines are conducted with their default hyper-parameter configurations.
+For a fair comparison, their image feature extractors are substituted with the ResNet101 (Khan et al., 2018).
+The training epoch is set to 10 and the batch size is set to 64, more than 10 will be overfitting, resulting in a decrease in accuracy.
+We take Adam as the optimizer weight decay 0.0005, and the learning rate is set to 0.0001.
+The $\epsilon$ in Eqn.
+3 is set to 0.9, the $\lambda_{1},...,\lambda_{4}$ in Eqn.
+3 are set to 10, 1, 10, and 1, respectively.
+To ensure the statistical significance of our experimental data, the average outcomes of each technique were calculated based on ten iterations.
+(a) PR Curve (AWA2)
+(b) PR Curve (CUB)
+(c) PR Curve (SUN)
+(d) P@N Curve (AWA2)
+(e) P@N Curve (CUB)
+(f) P@N Curve (SUN)
+(g) R@N Curve (AWA2)
+(h) R@N Curve (CUB)
+(i) R@N Curve (SUN)
+Figure 2. The comparison of PR Curve, P@N Curve and R@N Curve in 64 bit length codes.
+4.2. RQ1: Overall Performance
 (a) mAP@all in AWA2
@@
 (l) Decline Percentage (AUC) in SUN
-Figure 3. mAP@all and AUC scores and their decline percentage with the change of unseen classes ratio in the training process.
-3.1.1. Datasets
-In our investigations, we meticulously analyzed three distinguished zero-shot hashing datasets referenced in (Shi et al., 2022), i.e., AWA2 (Xian et al., 2018), CUB (Wah et al., 2011), SUN (Patterson and Hays, 2012).
-AWA2 encompasses 37,322 images across 50 animal categories, with a configuration of 40 seen and 10 unseen classes, each described by 85 attributes [39].
-The CUB dataset includes 11,788 images from 200 bird species, categorized into 150 seen and 50 unseen classes, each characterized by 312 attributes.
-SUN dataset, on the other hand, contains 14,340 images from 717 scene categories, with a division of 645 seen and 72 unseen classes, each depicted by 102 attributes.
-3.1.2. Baselines
-We select the following representative image hashing baselines as our competitors:
-  • Image Shallow Hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations.
-Typical works include LSH (Andoni and Indyk, 2008), SH (Weiss et al., 2008), ITQ (Gong et al., 2012), SDH (Shen et al., 2015a), IMH (Shen et al., 2015b), and PCA (Maćkiewicz and Ratajczak, 1993).
-Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
-  • Image Deep Hashing methods, leverage deep neural networks to automatically learn hierarchical and abstract representations of images.
-Representative works can be categorized into two paths, supervised hashing works, i.e, HashNet (Cao et al., 2017), GreedyHash (Su et al., 2018), JMLH (Shen et al., 2019), ADSH (Zhou et al., 2019), CSQ (Yuan et al., 2020), DPN (Fan et al., 2020), OrthoCos (Hoe et al., 2021), and unsupervised hashing works, i.e, BiHalf (Li and van Gemert, 2021), TBH (Qiu et al., 2021), CIBHash (Qiu et al., 2021).
-By employing end-to-end learning, deep hashing methods aim to optimize hash codes directly from raw pixel values, allowing for the discovery of intricate and discriminative image representations.
-  • Zero-shot Image Hashing methods extend the above algorithms to handle scenarios where the model encounters classes during testing that were not seen during training.
-These methods often incorporate semantic information, such as image attribute descriptions, or meta-learning strategies to enhance the model’s adaptability to unseen classes.
-Representative works include TSK (Yang et al., 2016), ASZH, SitNet, OPZH, and AH.
-ASZH (Shi et al., 2022) uses category labels to learn the hash functions from the seen training data, to generate binary codes for unseen data.
-SitNet (Guo et al., 2017), and OPZH (Zhang et al., 2019) adopt semantic vectors to find semantic embedding space for knowledge transferability.
-AH (Xu et al., 2017) constructs relationships and image features for transferability.
-3.1.3. Evaluation Metric
-For evaluating different aspects of our study, we employed a variety of metrics.
-These include the mean Average Precision (mAP) and the Area Under the Precision-Recall Curve (AUC) for assessing the ranking quality of search outcomes.
-Specifically, we present findings for the top 5,000 samples and for the entire dataset, denoted as mAP@5000 and mAP@all, respectively.
-Additionally, we utilized four conventional metrics: mean Average Precision (mAP), Precision-Recall Curve (PR Curve), Precision@N Curve (P@N Curve), and Recall@N Curve (R@N Curve), with a particular focus on the ranking performance for the initial 5,000 retrieved samples.
-3.1.4. Implementation Details
-All comparative baselines are conducted with their default hyper-parameter configurations.
-For a fair comparison, their image feature extractors are substituted with the ResNet101 (Khan et al., 2018).
-The training epoch is set to 10 and the batch size is set to 64, and more than 10 will be overfitting, resulting in a decrease in accuracy.
-We take Adam as the optimizer weight decay 0.0005 and the learning rate is set to 0.0001.
-The $\epsilon$ in Eqn.
-9 is set to 0.9, the $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ in Eqn.
-4 are set to 10, 1, and 10 respectively.
-To ensure the statistical significance of our experimental data, the average outcomes of each technique were calculated based on ten iterations.
-(a) PR Curve (AWA2)
-(b) PR Curve (CUB)
-(c) PR Curve (SUN)
-(d) P@N Curve (AWA2)
-(e) P@N Curve (CUB)
-(f) P@N Curve (SUN)
-(g) R@N Curve (AWA2)
-(h) R@N Curve (CUB)
-(i) R@N Curve (SUN)
-Figure 4. The comparison of PR Curve, P@N Curve and R@N Curve in 64 bit length codes.
-3.2. RQ1: Overall Performance
+Figure 3. mAP and AUC scores with the change of unseen classes ratio in the training process.
 Table.
 1 displays the results of the COMAE and all competitive baselines on datasets AWA2, CUB, and SUN with hash code lengths of 24, 48, 64, and 128.
-Note that the source codes of ASZH, SitNet, AH, and OPZH are unavailable, thus we use their results in paper (Shi et al., 2022).
+Note that the source codes of ASZH, SitNet, AH, and OPZH are unavailable, thus we use their results in the paper (Shi et al., 2022).
 The experimental settings of Table 1 are the same as (Shi et al., 2022).
 In Fig.
-4, we illustrate three distinct curves, i.e., PR Curve, R@N Curve, and the P@N Curve, to evaluate the performance of hash codes with a length of 64 bits on all datasets.
+2, we illustrate three distinct curves, i.e., PR Curve, R@N Curve, and the P@N Curve, to evaluate the performance of hash codes with a length of 64 bits on all datasets.
 Typically, the efficacy of a method is gauged by the PR curve approaching the upper right quadrant, with higher positions on the P@N and R@N curves indicative of superior performance.
-Overall, we find the following conclusions,
-  • As shown in Table.
+We find the following conclusions: (1) As shown in Table.
 1, deep hashing models perform better than shallow hashing ones in general, which shows the superiority of neural networks.
 In addition, the zero-shot hashing models exhibit superior performance in comparison to both deep and shallow hashing models.
-  • COMAE surpasses all existing zero-shot hashing benchmarks, in terms of all bit lengths.
+(2) COMAE surpasses all existing zero-shot hashing benchmarks, in terms of all bit lengths.
 At the same time, the performance of COMAE remains relatively stable when the number of bits increases.
-Compared with AH (Xu et al., 2017), an attribute-based zero-shot hashing method, we gain better performance, which serves to validate the correctness and effectiveness of the motivation we put forth.
-  • An insightful observation can be derived from the graphical representation in Fig.
-4 in all datasets.
-COMAE attains a heightened accuracy at a comparatively lower recall rate, thereby demonstrating the superiority of our proposed method.
-3.3. RQ1: Hashing Code Visualization
-In this experiment, we adopted a more intuitive approach to visualizing hashing codes to evaluate the effectiveness of our proposed COMAE.
-Throughout the comparative experiments, we initially transformed the 64-bit hashing codes into lower-dimensional spaces, subsequently projecting these codes into a 2-D plane using the t-SNE (Van der Maaten and Hinton, 2008) tool.
-This methodology facilitated a clearer observation of COMAE’s performance through visual means.
-We present the visualization of hash code for all testing samples with unseen classes on the AWA2 dataset.
-Results are shown in Fig.
-7.
-According to the visualization results, we find that a) HashNet, DPN and OrthoCos perform badly, which are unable to distinguish the unseen classes.
-b) Although CSQ and BiHalf gain favorable outcomes in visualization and mAP@5000 results in Table.
-1, they still encounter challenges such as substantial inter-class overlap and indistinct boundaries (the classes represented by the green and other colors).
-c) Compared with other baselines, the proposed COMAE achieves the best performance, as the generated hash codes are more distinguished.
-It indicates that COMAE can effectively classify images.
+Compared with PIXEL, the latest attribute-based zero-shot hashing method, we gain better performance, which serves to validate the correctness and effectiveness of the motivation we put forth.
+(3) Based on the results shown in Fig.
+2 across all datasets, COMAE attains a heightened accuracy at a comparatively lower recall rate, thereby demonstrating its superiority.
 (a) without $\mathcal{L}_{\text{classwise}}$
@@
 4.3. RQ2: Effects of Unseen Class Ratio
-To verify the robustness of our model in the context of the zero-shot scenario, we conduct experiments where the number of seen class categories underwent a progressive reduction during the training process.
+To verify the robustness of our model in the context of the zero-shot scenario, we conduct experiments where the number of seen class categories undergoes a progressive reduction during the training process.
 The results are visually depicted in Fig.
@@
 Specifically, (d)-(f) and (j)-(l) show the performance decrease ratio between 20%-¿40% and 20%-¿60% respectively.
-We conclude the following conclusions based on the outcome results,
-  • All deep hashing methods manifest a discernible decline in performance concomitant with an escalation in the number of unseen classes on three datasets.
-Notably, this performance deterioration not only intensifies but also amplifies in magnitude, as evidenced by both the MAP@all and AUC metrics.
-  • COMAE exhibits a performance superiority across various unseen class ratios, along with the most minimal decrease.
-(a) $\lambda_{1}$
-(b) $\lambda_{2}$
-(c) $\lambda_{3}$
-(d) $\epsilon$
-Figure 6. The effects of hyper-parameters on AWA2 dataset.
-(a) COMAE
-(b) HashNet
-(c) DPN
-(d) OrthoCos
-(e) CSQ
-(f) BiHalf
-Figure 7. t-SNE visualizations of hashing code on test unseen classes. The 10 colors denote 10 different unseen classes.
-3.5. RQ3: Distance Visualization of Intra and Inter Classes
-To verify whether the classwise constraint can broaden the distance of inter-class and reduce the distance of intra-class, we provide a quantitative comparison of the distance between the mean value of intra-class distances and inter-class distances, under the setting of without loss $\mathcal{L}_{\text{contrastive}}$ and with the loss.
+Based on the outcome results, we conclude the following conclusions: (1) All deep hashing methods show a noticeable decline in performance along with an increase in the number of unseen classes on three datasets.
+Notably, this performance drop not only intensifies but also amplifies in magnitude, as evidenced by both the MAP@all and AUC metrics.
+(2) COMAE exhibits a performance superiority across various unseen class ratios, with minimal decrease.
+Figure 5. Case study on the CUB dataset.
+Table 2. The average running time for each epoch
+Method AWA2 CUB SUN
+mAP Train Infer. mAP Train Infer. mAP Train Infer.
+PIXEL 0.5456 483s 0.154s 0.2819 625s 0.183s 0.1741 568s 0.231s
+COMAE 0.5801 45s 0.011s 0.2532 48s 0.052s 0.1764 47s 0.075s
+4.4. RQ3: Case Study
+We illustrate the performance of COMAE through case studies with 64-bit hash codes, as shown in Fig 5.
+We present the top 10 retrieved images for each query using COMAE (first row) and PIXEL (second row).
+Correctly retrieved images are outlined in green, while incorrect ones are outlined in red.
+COMAE demonstrates superior retrieval performance, effectively capturing key visual features.
+For queries with distinctive visual characteristics, such as the yellow-headed blackbird or the blue bunting, COMAE retrieves highly accurate results, primarily by leveraging attributes like feather color, body shape, and posture.
+In more challenging cases involving birds with subtle differences or overlapping features, such as cuckoos and flycatchers, COMAE still maintains strong retrieval accuracy but occasionally retrieves images with semantically similar attributes.
+For instance, brownish birds with similar body structures sometimes lead to false positives, as observed in some incorrect retrievals.
+Overall, for intact retrieval targets, COMAE achieves an accuracy of 80%, while in complex cases involving partial occlusion, highly similar species, or ambiguous visual features, COMAE retains a 70% accuracy rate, outperforming PIXEL in zero-shot hashing tasks.
+The comparison between COMAE and PIXEL highlights COMAE’s robustness in learning effective feature representations and reducing semantic drift in image retrieval.
+4.5. RQ4: Efficiency Analysis
+We conduct model efficiency analysis with a 64-bit hashing length in this section.
+Table 2 reports the average training time per epoch and inference time per epoch.
+We select the best and latest baseline, PIXEL, as our comparable baseline.
+We find that COMAE achieves competitive computational efficiency.
+In the largest dataset, SUN, COMAE reduces the training time by over 91.7%, and the inference time is reduced from 0.231s (PIXEL) to 0.075s (COMAE), with an improvement of 67.5%.
+This demonstrates that COMAE not only trains significantly faster but also achieves efficient inference, making it well-suited for real world large-scale retrieval systems.
+4.6. RQ5: Model Analysis
+Distance Visualization of Intra and Inter Classes.
+For model analysis, we first conduct the distance visualization of intra and inter classes, to verify whether the classwise constraint can broaden the distance of inter-class and reduce the distance of intra-class.
+We provide a quantitative comparison of the distance between the mean value of intra-class distances and inter-class distances, under the setting of without the contrastive loss $\mathcal{L}_{\text{pairwise}}$ and with it.
 Fig.
 4 summarizes the histogram of intra-class and inter-class distances on the AWA2 dataset.
-We adjust the frequency normalization so the total of all histogram bins equals one.
+We adjust the frequency normalization so that the total of all histogram bins equals one.
 The directional marker indicates the separability in Hamming distances, $\mathbb{E}[D_{inter}]-\mathbb{E}[D_{intra}]$.
@@
 The results show that our proposed classwise constraint can enlarge the distance of the inter-class and reduce the distance of the intra-class.
-3.6. RQ3: Component Analysis
+Component Analysis.
 To provide further insight into COMAE, we conduct component analysis on AWA2 and CUB datasets with 64-bit hashing-code lengths.
-We configure several variants of COMAE as a) COMAE-V1 is the COMAE model without $\mathcal{L}_{\text{pointwise}}$; b) COMAE-V2 represents the COMAE without $\mathcal{L}_{\text{pairwise}}$; c) COMAE-V3 refers to the COMAE without $\mathcal{L}_{\text{classwise}}$.
+We configure several variants of COMAE as (1) COMAE-V1 is the COMAE model without $\mathcal{L}_{\text{pointwise}}$; (2) COMAE-V2 represents the COMAE without $\mathcal{L}_{\text{pairwise}}$; (3) COMAE-V3 refers to the COMAE without $\mathcal{L}_{\text{classwise}}$.
 Results are shown in Table 3.
 Compared with other components, the pointwise objective plays the most significant role in AWA2, and the pairwise objective in CUB.
-Also, the performance of COMAE on both datasets surpasses that of variants, validating the necessity of each component.
-Table 2. The mAP@all and AUC results with switching model components in different datasets.
-AWA2 CUB
-mAP@all AUC mAP@all AUC
-COMAE-V1 0.4830 0.4781 0.2069 0.2418
-COMAE-V2 0.4857 0.4791 0.1867 0.2240
-COMAE-V3 0.4835 0.4782 0.2064 0.2412
-COMAE 0.4883 0.4812 0.2094 0.2416
-3.7. RQ4: Parameter Analysis
-We analyze the impact of critical hyper-parameters, i.e., $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ that influence the weight of the auxiliary losses in Eqn.
-9 and the threshold $\epsilon$ for selecting positive samples in Eqn.
-4.
-These evaluations were performed using the AWA2 dataset with 64-bit hash codes, and the results are detailed in Fig.
-6.
-As the increase of hyper-parameters, the performance of our model initially rises and then decreases.
-Hence, we choose the best hyper-parameters corresponding to the point at peak performance.
+Also, the performance of COMAE on both datasets surpasses that of the variants, validating the necessity of each component.
+Table 3. The mAP@5000 and AUC results with switching model components in different datasets.
+AWA2 CUB SUN
+mAP@5000 AUC mAP@5000 AUC mAP@5000 AUC
+COMAE-V1 0.5093 0.4364 0.1988 0.2333 0.1420 0.1286
+COMAE-V2 0.5122 0.4981 0.1794 0.2161 0.1310 0.1243
+COMAE-V3 0.5099 0.5025 0.1983 0.2327 0.1350 0.1264
+COMAE 0.5149 0.5106 0.2012 0.2331 0.1465 0.1302
 5. Conclusion
+Zero-Shot Hashing stands out for its high efficiency and robust generalization, particularly in large-scale retrieval scenarios.
 In this paper, we present COMAE for a comprehensive attribute exploration in ZSH.
 COMAE depicts the relationships between seeable and unseeable classes through meticulously designed explorations: point-wise, pair-wise, and class-wise constraints.
+Furthermore, theoretical analysis is provided to show the effectiveness of COMAE.
 Our model employs an attribute prototype network to regress attributes, facilitating the learning of local features.
@@
 Finally, we introduce the class-wise constraint to cohesively enhance hashing learning, image representations, and visual attributes.
-Experiments demonstrate the excellent performance of COMAE.
+Experimental results on the popular ZSH datasets demonstrate that COMAE outperforms state-of-the-art hashing techniques, especially in scenarios with a larger number of unseen label classes.
+Acknowledgements.
+We would like to thank Zeyu Dong, Ning Cao, Shuai Liu, Pengfei Wang, and Yuchen Yan who discussed with us and provided instructive suggestions.
+This work is partially supported by the National Natural Science Foundation of China (No.92470204), the Beijing Natural Science Foundation (No.4254089), the Postdoctoral Fellowship Program of CPSF (No.GZC20232736), and the China Postdoctoral Science Foundation Funded Project (No.2023M743565).